the resultant of two perpendicular forces has a magnitude of 40 n. if the magnitude of one of the forces is 25 n, what is the magnitude of the other force?


1
Expert's answer
2021-05-03T10:28:20-0400
"f=\sqrt{f_x^2+f_y^2}\\40=\sqrt{25^2+f_y^2}\\f_y=31\ N"